OpenStudy (anonymous):

Use the right hand endpoint of the Riemann sum. f(x)=cosx on [0, pi]; n=6

OpenStudy (anonymous):

\[\Delta x=(b-a)/n = (\pi-0)/6 = \pi/6\] Now: \[\S = \Delta x (f(0) + f(0+\pi/6)+f(0+2*\pi/6) + ... + f(b))\] Simply compute...

OpenStudy (nikvist):

right hand endpoint \[S=\Delta x\left(f(\pi/6)+f(2\pi/6)+f(3\pi/6)+f(4\pi/6)+f(5\pi/6)+f(\pi)\right)\]

OpenStudy (anonymous):

Ops, nik is correct, essentially the same as mine but I added f(0) by mistake.